# AgentMove: Predicting Human Mobility Anywhere Using Large Language Model based Agentic Framework

### v1

@@ Methods
We use LibCity (Jiang et al. 2023) to implement the FPMC, RNN, DeepMove and LSTPM. We follow the default parameter settings of these models in the library for training and inference. For LLMs, we use OpenAI API  [^4] https://platform.openai.com/ for accessing GPT4omini and DeepInfra  [^5] https://deepinfra.com/ for accessing other open source LLMs, including Llama3, Qwen2, Gemma2 and GLM4. Detailed parameter settings for those baselines can be found in the appendix.
Main Results
In this section, we compare AgentMove with 6 baselines in 4 cities to demonstrate the effectiveness of proposed framework. Here, we use GPT4omini as the default base LLM of AgentMove and the results of AgentMove with other kinds of base LLMs can be found in first section A in appendix. To ensure a fair comparison, the LLM-based baselines, LLM-Mob and LLM-ZS, also employ the same base LLM as AgentMove. The results are presented in Table 1.
@@
As the representative deep learning models, DeepMove and LSTPM achieve best or second-best results in 5 out of 12 metrics, and both of their performance varies among different cities. Compared with the deep learning baselines, the best LLM-based baseline LLM-ZS can achieve better results than DeepMove and LSTPM in 7 out of 12 metrics. The promising results of LLM-ZS present the powerful sequential pattern discovery and reasoning ability of LLM in modeling mobility, which demonstrate the potential of the LLM-based method. It is noted that the results of LLM-based methods are zero-shot prediction while the deep learning based methods rely on sufficient training with enough mobility data. Compared with these baselines, our proposed method AgentMove is the best method and achieves the best results in 11 out of 12 metrics in 4 datasets. Compared with the deep learning methods, AgentMove outperforms them more than 20% in half of the metrics. Even in the weakest results, it performs very close to the strongest baseline DeepMove with less than 3% performance drop. Compared with the advanced LLM-based baselines, AgentMove outperforms both of them more than 5.7%-39.4% in all the metrics. These results demonstrate the effectiveness of proposed framework in stimulating the comprehensive ability of LLM-base agentic framework for mobility prediction.
Ablation Study
(a) The length of the input history of memory unit.
(b) The number of generated places in spatial world.
Figure 4: Effects of two key parameters in AgentMove.
In this section, we conduct ablation study to demonstrate the contribution of each component in AgentMove. Related results are presented in Table 2 and Figure 4. Similar to the last section, we use GPT4omini as the base LLM of AgentMove.
@@
We first discuss the impact of three core components individually, as detailed in the top four lines of Table 2. We have two important observations: 1) each core component not always works well for all the datasets, they may hurt the performance in some cases. For example, after adding spatial temporal memory design in the base prompt, the performance of AgentMove in Tokyo is decreased significantly, but it performs much better in Moscow, 2) the performance gain of each component in different metrics are also different. For example, while memory design leads to better performance in the Acc@1 in SaoPaulo, the performance in other three metrics are dropped. Then we discuss the effects of the combination of the core components in the last three lines in Table 2. We find that while memory or world model generator can not improve the prediction performance individually, their combination introduces significant performance gain in all the metrics of Tokyo dataset. After adding the collective knowledge extractor, the performance gain becomes larger in most of the metrics of Tokyo and SaoPaulo. In summary, compared with the base prompt design, the combination of proposed designs introduce 7%-45% performance gain in all the datasets.
In Figure 4, we further discuss the effects of several important parameters in the core component of AgentMove, including the history length in memory unit, the number of potential locations generated in the word knowledge generator. When studying the effects of one parameter, we keep all the left parameters fixed. As Figure 4(a) shows, the performance of AgentMove generally improves as the length of historical input in the memory unit increase. Figure 4(b) presents the effects of the number of potential locations based on the associations with urban spatial structures, which is varying a lot among different cities.
Models FSQ@Tokyo FSQ@SaoPaulo FSQ@Moscow ISP@Shanghai
base 0.175 0.410 0.299 0.165 0.385 0.277 0.120 0.340 0.233 0.170 0.425 0.298
+memory 0.145 0.395 0.274 0.190 0.315 0.255 0.160 0.365 0.265 0.170 0.445 0.312
+world 0.150 0.395 0.280 0.175 0.365 0.269 0.090 0.335 0.213 0.155 0.390 0.276
+collective 0.150 0.435 0.298 0.175 0.380 0.275 0.140 0.340 0.246 0.175 0.465 0.317
+mem+world 0.195 0.455 0.328 0.240 0.390 0.317 0.135 0.365 0.254 0.215 0.455 0.342
+mem+world+col 0.185 0.465 0.331 0.230 0.415 0.326 0.150 0.360 0.259 0.190 0.450 0.329
vs base 11.43% 13.41% 10.98% 45.45% 7.79% 17.99% 33.33% 7.35% 13.91% 11.76% 5.88% 10.30%
@@
Table 2: Ablation study results of AgentMove. ‘base’ denotes the basic prompts which is similar to the baseline LLM-ZS, ‘+memory’ or ‘+mem’ denotes adding spatial-temporal memory, ‘+world’ denotes adding world knowledge generator, ‘+collective’ or ‘+col’ denotes adding collective knowledge extractor.
Geographical Bias Analysis
@@
Figure 5: Geospatial bias analysis of 3 LLM-based methods.
While LLMs are trained with the online web text which can be geographically bias (Manvi et al. 2024) around the world. We investigate the potential geographical bias in LLM based mobility prediction methods and attempt to answer whether AgentMove can alleviate the geographical bias inherent in LLMs to some extent. Experiment results conducted on 12 cities are presented in Figure 5.
In Figure 5(a), we can find significant differences in the accuracy of three LMM-based methods across cities. For instance, cities like Tokyo, Paris, and Sydney generally achieve higher accuracy, while cities like Cape Town and Nairobi see notably lower performance. This suggests the presence of geographical bias in trained LLMs. We also find that proposed AgentMove performs best in most of the cities. Figure 5(b) provides a box-plot test by comparing the Acc@5 of the three LLM-based methods in 12 cities. Results demonstrate that AgentMove not only outperforms the other methods in terms of overall accuracy but also exhibits a smaller range of error. The performance of AgentMove is more consistent across different cities, suggesting a reduced impact of geographical bias.
Related Work
Mobility Prediction with Deep Learning
Significant efforts have been made in mobility prediction using deep learning models, encompassing research from both sequential-based methods and graph-based methods. While traditional methods typically employ Markov models (Rendle, Freudenthaler, and Schmidt-Thieme 2010; Cheng et al. 2013) to predict the next visit by learning the transition probabilities between consecutive POIs, sequential-based deep learning methods are first proposed to model the high-order movement patterns of trajectory data. It contains two kinds of works, Recurrent Neural Networks (RNNs) (Kong and Wu 2018; Huang et al. 2019; Yang et al. 2020; Zhao et al. 2020), and attention mechanisms(Feng et al. 2018; Luo, Liu, and Liu 2021; Lin et al. 2021; Cui et al. 2021; Qin et al. 2022; Hong et al. 2023) based works. Despite their success, these methods primary focus on extracting mobility patterns from an individual perspective, while overlooking the collaborative information available from other users’ trajectories. To this end, recent works (Rao et al. 2022; Yang, Liu, and Zhao 2022) have explored the use of Graph Neural Networks (GNNs) for their ability to model complex relationships. However, all these methods rely on collecting large volume of private trajectory. Our proposed AgentMove leverages the world knowledge and sequential modeling abilities of LLM to enable the generalized mobility prediction with zero-shot prediction ability.
Large Language Models and Agents
Due to the powerful language based generalization and reasoning capabilities (Wei et al. 2022a), large language models (OpenAI 2022; Touvron et al. 2023) develop rapidly and have been widely applied in various tasks, e.g., coding (Qian et al. 2024) and math (Wei et al. 2022a). Recent works (Gurnee and Tegmark 2023; Manvi et al. 2023) find that LLMs have learned lots of geographical knowledge of the world. Besides, researcher also explore the potential of applying LLMs in spatial temporal data modelling by directly transferring the domain specific task into the language format, e.g., time series forecasting (Gruver et al. 2024), traffic prediction (Li et al. 2024b), trajectory mining (Wang et al. 2023; Beneduce, Lepri, and Luca 2024), trip recommendation (Xie et al. 2024; Li et al. 2024a) and so on. These early works present the potential of applying LLMs in spatial temporal modelling. To effectively utilize the vast knowledge hidden in the LLMs and stimulate their reasoning and planning abilities, lots of prompts techniques (Wei et al. 2022b; Kojima et al. 2022; Wang et al. 2022; Yao et al. 2024) are proposed in solving naive text games and math problem. However, when it comes to the complicated task in the real life and domain-specific tasks, only the prompt techniques are not enough. Recently, LLM based agents (Wang et al. 2024; Xi et al. 2023) are proposed to solve this question by equipping LLMs with explicit memory, working flow and external tools. Here, we are the first to design LLM based agent for mobility prediction task. With explicit spatial temporal memory and work flow of spatial structure and social structure mining, we succeed in leveraging the world knowledge of LLM and structured reasoning capability for spatial temporal trajectory modelling.
Conclusion
In this paper, we propose AgentMove, a systematic agentic framework for evolvable and generalized mobility prediction around the world. In AgentMove, we first decompose the mobility prediction task into three sub-tasks. To solve these sub-tasks, we design spatial-temporal memory and collective knowledge extractor to learn the individual mobility pattern and shared mobility pattern among users. Furthermore, we design a world knowledge generator to make use of the text address to understand the urban structures like human and thus enable the effective exploration modelling of human mobility. Extensive experiments on mobility trajectories from 12 cities demonstrate the superiority and robustness of AgentMove. In the future, we plan to further explore the better ways to extract and utilize the vast world knowledge and common sense of LLM for mobility modelling. Besides, we will explore the framework in other mobility trajectory related tasks, e.g., trajectory generation.
Appendix A Appendix
The Impact of Different LLMs
As the core foundation of AgentMove, the strength of base LLM’s capabilities has a significant impact on the performance of AgentMove. Thus, we evaluate the impact of different LLMs with varying sizes and sources in Figure 6. We consider the influence of LLM size, LLM sources on the performance of three LLM based methods in Figure 6(a) and Figure 6(c). We dive into the detailed impacts of the LLM size and sources on AgentMove among three cities in Figure 6(b) and Figure 6(d).
We first investigate how the LMM size affects the performance of LMM-based methods. In Figure 6(a), we find that LLMs with more parameters, such as Llama3-70B, generally outperform smaller model like LLama3-8B in the LLM-based methods, highlighting the advantage of increased model capacity in helping mobility prediction. It is also noted that the performance gain from scaling up the Llama3 is very small. The only exception is that LLM-Mob in LLama3.1-405B performs much worse. After checking the results carefully, we find that this is caused by the wrong output format of LLM-Mob@Llama3.1-405B with too much detailed reasoning steps, which demonstrates the potential instability of the LLM-Mob among various LLMs. In Figure 6(c), we shift our focus to the impact of various 7B LLM with different training data and model structures. The results show that proposed AgentMove performs best adaptability among different LLMs. While LLM-Mob performs stable in all the 7B-LLMs, its performance on Gemma2-9B is far worse than other two methods.
We then discuss the detailed impacts of LLM size and source specifically on AgentMove’s performance across three cities. Figure 6(b) reveals that that larger models, particularly Llama3.1-405B, generally deliver significant performance gains for AgentMove compared to smaller models like Llama3-8B across different cities. It is also observed that in Tokyo, Llama3-1-405B performs slightly weaker compared to Llama3-70B. This suggests that while larger models often excel, their effectiveness may vary depending on the unique mobility patterns and characteristics of each city. Figure 6(d) further shows the role of LMM with different training data and model structures in determining AgentMove’s performance at different cities. Although the source Llama3-1-405B often leads in accuracy, other kinds of LLM sometimes perform similarly or better in certain contexts. In general, considering the cost and robustness, GPT4omini is a good start point for utilizing AgentMove in the mobility prediction task.
(a) Effects of LLM size on the performance of three LLM based methods.
(b) Effects of LLM size on the performance of AgentMove for three cities.
(c) Effects of LLM from different source on the performance of three LLM-based methods
(d) Effects of LLM from different source on the performance of AgentMove for three cities
@@ Appendix A Appendix
Figure 6: The effects of LLM with varying sizes and sources on the prediction performance of three LLM based methods. Only results from Accuracy@5 are presented here, when similar trends are observed on other metrics.
Effects of Different Location Representation
We explore the effects of the format of location ID in the prompt in Figure 7, where ‘int’ denotes to re-encoding the locations into ordered number in each city, ‘str’ denotes preserve the original string ID of venue in Foursquare. We find that in all the tree datasets, AgentMove with ‘int’ encoding strategy achieves the similar performance in two metrics compared with the representation of ‘str’ location ID. This result also demonstrate that the outstanding prediction performance of LLM based methods does not stem from the potential data leakage. Thus, all the results in our experiments use the ‘str’ encoding as the default encoding strategy.
Figure 7: Effects of the format of location ID, where ‘str’ denotes the original location string in Foursquare, ‘int’ denotes that encoding the location into ordered integer in each city.
Other results of Geospatial Bias
Following the results in section Geographical Bias Analysis, we reports the performance of AgentMove variants, as illustrated in Table 2, among different cities in Figure 8(a). In most cases, the performance fluctuation between different AgentMove variants in different cities are close. We also provide a analysis by comparing the Acc@1 of the three LLM-based methods at different cities in Figure 8(b).
(a) Performance of different variants of AgentMove on 12 cities.
(b) The distribution of Acc@1 across three methods on 12 cities.
Figure 8: Geospatial bias analysis of three LLM-based methods on the mobility prediction.
Details of Data
Detailed information of processed trajectory data from 12 cities is presented in Table 3.
City Users Trajectories Locations Days Records
Tokyo 12464 112942 83190 531 1030105
SaoPaulo 11856 77120 78904 531 809198
Moscow 10501 100854 93599 531 950898
NewYork 15785 28502 41386 531 380247
Sydney 1720 4557 10523 531 54250
Paris 6903 7559 19837 531 111325
London 9724 14596 28687 531 188530
Beijing 1076 1847 5753 531 21030
Shanghai-FSQ 1272 3238 8014 531 33129
Shanghai-ISP 1762 2844 12576 7 325215
Capetown 403 1234 2988 531 13303
Mumbai 1070 3070 7942 531 40592
Nairobi 356 2690 5807 531 28453
Table 3: Trajectory statistics of 12 cities around the world.
Prompt Examples
Here, we present the detailed prompts for each LLM based methods.
Prompt of AgentMove
1## Task
2Your task is to predict <next_place_id> in <target_stay>, a location with an unknown ID, while temporal data is available.
3
4## Predict <next_place_id> by considering:
51. The users activity trends gleaned from <historical_stays> and the current activities from <context_stays>.
62. Temporal details (start_time and day_of_week) of the target stay, crucial for understanding activity variations.
73. The potential places that users may visit based on an overall analysis of multi-level urban spaces.
84. The personal profile and memory info extracted from the long trajectory history of each user.
9
10## The potential places from the global spatial view:
11{spatial_world_info}
12
13## The nearby places visited by other users with similar mobility pattern:
14{social_world_info}
15
16## The personal profile and long memory:
17{spatial_temporal_memory_info}
18
19## The history data:
20<historical_stays>: {historical_stays}
21<context_stays>: {context_stays}
22<target_stay>: {target_time, <next_place_id>}
23
24## Output
25Present your answer in a JSON object with:
26"prediction" (list of IDs of the five most probable places, ranked by probability) and "reason" (a concise justification for your prediction).
Prompt for spatial-temporal memory unit.
1### long term memory info
2Place id to name mapping: {venue_id_to_name}.
3In historical stays, The user frequently engages in activities at {frequent_hours}.
4The most frequently visited venues are {frequent_venues}.
5Hourly venue activities include {hourly_activity_desc}.
6The users activity transitions often include sequences such as {transitions}.
7
8### short term memory info
9In recent context stays, users last visit was on {}
10Frequently visited locations include: {}
11Visit times: {}
12
13### user profile
14The user is most active at {most_frequent_hour} with {most_frequent_count} visits.
15They frequently visit {most_frequent_venue_category} with {most_frequent_venue_count} visits
16Based on the data, the user {’, ’.join(insights)}.
Prompts for world knowledge generator.
1# Prompts for world knowledge generator
2
3## prompts for extracting structured address info
4{original address info from https://nominatim.org/ by querying via reverse API}
5Please get the administrative area name, subdistrict name/neighbourhood name, access road or feeder road name, building name/POI name.
6Present your answer in a JSON object with:’administrative (the administrative area name) ,’subdistrict (subdistrict name/neighbourhood name),’poi’(building name/POI name),’street’(access road or feeder road name which POI/building is on).
7Do not include the key if information is not given.Do not output other content.
8
9### block info
10This trajectory moves within following administrative areas:
11{administrative_area}
12This trajectory sequentially visited following subdistricts, with the last subdistrict being the most recently visited:{}
13Consider about following two aspects:
141.The frequency each subdistrict is visited.
152.Transition probability between two administrative areas.
16Please predict the next subdistrict in the trajectory. Give {explore_num} subdistricts that are relatively likely to be visited. Do not output other content.
17
18### poi and street info
19This trajectory sequentially visited following POIs(Each POI is represented by POI name, the feeder road or access road it is on’), with the last POI being the most recently visited:{pois})
20Consider about following two aspects:
211.The frequency each subdistrict is visited.
222.The frequency each poi is visited.
233.Transition probability between two subdistricts.
244.Transition probability between two pois.
25Please predict the next poi in the trajectory.Give {explore_num} POIs that are relatively likely to be visited. Do not output other content.
26
27
28# spatial world model info used in AgentMove
29### Names of subdistricts that are relatively likely to be visited:
30{block_info}
31### Names of POIs that are relatively likely to be visited:
32{poi_info}
Prompt for collective knowledge extractor.
1## Finding neighbors
2neighbors = list(graph.neighbors(venue_id))
3sorted_neighbors_freq = [(n, 1) for n in neighbors if n not in context_trajs]
4
5## Prompts in final reasoning step
61-hop neighbor places in the social world: {neighbors}
7......
Prompt of LLM-Mob
1Your task is to predict a users next location based on his/her activity pattern.
2You will be provided with <history> which is a list containing this users historical stays, then <context> which provide contextual information
3about where and when this user has been to recently. Stays in both <history> and <context> are in chronological order.
4Each stay takes on such form as (start_time, day_of_week, duration, place_id). The detailed explanation of each element is as follows:
5start_time: the start time of the stay in 12h clock format.
6day_of_week: indicating the day of the week.
7duration: an integer indicating the duration (in minute) of each stay. Note that this will be None in the <target_stay> introduced later.
8place_id: an integer representing the unique place ID, which indicates where the stay is.
9
10Then you need to do next location prediction on <target_stay> which is the prediction target with unknown place ID denoted as <next_place_id> and
11unknown duration denoted as None, while temporal information is provided.
12
13Please infer what the <next_place_id> might be (please output the 10 most likely places which are ranked in descending order in terms of probability), considering the following aspects:
141. the activity pattern of this user that you learned from <history>, e.g., repeated visits to certain places during certain times;
152. the context stays in <context>, which provide more recent activities of this user;
163. the temporal information (i.e., start_time and day_of_week) of target stay, which is important because peoples activity varies during different time (e.g., nighttime versus daytime)
17and on different days (e.g., weekday versus weekend).
18
19Please organize your answer in a JSON object containing following keys:
20"prediction" (the ID of the five most probable places in descending order of probability) and "reason" (a concise explanation that supports your prediction). Do not include line breaks in your output.
21
22The data are as follows:
23<historical>: {historical_stays}
24<context>: {context_stays}
25<target_stay>: {target_time, <next_place_id>}
Prompt of LLM-ZS
1Your task is to predict <next_place_id> in <target_stay>, a location with an unknown ID, while temporal data is available.
2
3Predict <next_place_id> by considering:
41. The users activity trends gleaned from <historical_stays> and the current activities from <context_stays>.
52. Temporal details (start_time and day_of_week) of the target stay, crucial for understanding activity variations.
6
7Present your answer in a JSON object with:
8"prediction" (IDs of the five most probable places, ranked by probability) and "reason" (a concise justification for your prediction).
9
10The data:
11<historical_stays>: {historical_stays}
12<context_stays>: {context_stays}
13<target_stay>: {target_time, <next_place_id>}
Parameter settings
Detailed parameter settings for each Markov and deep learning based baselines are presented in Table 4. For each baseline, we adapt the early stopping methods by considering the accuracy of validation set and learning rate schedule threshold. All the experiments of deep learning baselines are running on a machine with 64 cores, 512GB of memory, and 2 NVIDIA RTX 4090 GPU, which is installed with Ubuntu 22.04.3 LTS.
Table 4: Detailed parameter settings for Markov and deep learning based baselines.
Parameters FPMC RNN DeepMove LSTPM
batch size 1024 1024 128 128
learning rate (lr) - 1e-3 1e-3 1e-3
lr schedule step - 2 3 2
lr schedule decay - 0.1 0.1 0.1
schedule threshold - 1e-3 1e-3 1e-3
early stop lr - 9e-6 9e-6 9e-6
L2 - 1e-5 1e-5 1e-6
max epoch 100 30 30 30
loc embed size 64 500 500 500
hidden embed size - 500 500 500
dropout - 0.3 0.5 0.8
All the generation parameter settings for LLM based methods are the same. The temperature is set as 0 for deterministic results, the maximum output token is 1000, the maximum input token is 2000, other parameters are not set and follow the default settings from API provider.
Data preprocessing
Preprocessing for Foursquare Data
As introduced in section Datasets, we select 12 cities around the world to evaluate the performance of proposed framework. We match each trajectories with the target cities by calculating the minimum distance to the city center. For the ordered trajectories in each city, we use 72 hours as the time window to split the trajectory into sessions. We filter the users with less than 5 sessions and filter sessions with less than 4 stays. Then, we divide each trajectory dataset into training, validation, and test sets in a ratio of 7:1:2. During the testing, we filter the users with less than 3 sessions or more than 50 sessions which is designed to ensure the quality of testing users and also balance the effects from different users. Different from the previous works, we do not specifically filter locations. All the users and trajectories of them are sorted by the id. We select one session of each user and aggregate the first $n$ sessions from all the users to calculate the average accuracy. Here, $n$ is utilized to control the cost of evaluation for LLMs and keep fixed in the experiment, which is set as 200 in most of the experiments. It is noted that only the volume of testing set is controlled for cost, the entire training set is provided to the deep learning based methods for training.
Preprocessing for ISP Data
Following the preprocessing in the original paper (Feng et al. 2019), we split the data into different sessions by merging trajectory points in the same day. Due to the regularity of human, there are too much repeated trajectory points in the original sessions. To make the prediction challenging, we compress the trajectory sessions by merging the same locations within a time window (2 hours) and ignoring the visiting occurred during the night (from 8 p.m. to 8 a.m.). While the ISP data lasts only 7 days, we split the whole data into training set, validation set and testing data in a ratio of 4:1:5 for preserving enough testing data. The minimum session filter parameter is changed from 3 to 1.

### v2

@@ 4.1.3 Implementation
We use LibCity Jiang et al. (2023) to implement the FPMC, RNN, DeepMove and LSTPM. We use the official codes from author to implement GETNext  [^4] https://github.com/songyangme/GETNext and STHGCN [^5] https://github.com/ant-research/Spatio-Temporal-Hypergraph-Model. We follow the default parameter settings of these models in the library and official codes for training and inference. For LLMs, we use OpenAI API  [^6] https://platform.openai.com/ for accessing GPT4omini, DeepInfra  [^7] https://deepinfra.com/ and SiliconFlow  [^8] https://siliconflow.cn/models for accessing other open source LLMs. Detailed parameter settings for those baselines can be found in the appendix.
4.2 Main Results
In this section, we compare AgentMove with 9 baselines in 4 cities at Table 1. In the experiments, we use GPT4omini as the default base LLM for all LLM-based methods.
@@ 4.2 Main Results
As the representative deep learning models, GETNEext and STHGCN achieve best or second-best results in 4 out of 12 metrics. Compared with the deep learning baselines, the best LLM-based baseline LLM-Move can achieve better results than GETNext and STHGCN in 3 out of 12 metrics, which present the powerful sequential pattern discovery and reasoning ability of LLM in modeling mobility. It is noted that the results of LLM-based methods are zero-shot prediction while the deep learning based methods rely on sufficient training with enough mobility data. Compared with these baselines, our proposed method AgentMove is the best method and achieves the best results in 8 out of 12 metrics in 4 datasets. Although AgentMove falls slightly behind the best baseline, GETNext, in three metrics, two of them are very close. These results in Table 1 demonstrate the effectiveness of proposed framework in stimulating the comprehensive ability of LLM-based agentic framework for mobility prediction.
4.3 Ablation Study on Model Designs
In this section, we provide a more detailed analysis of the proposed method under varying model designs to further demonstrate its effectiveness.
@@ 4.3 Ablation Study on Model Designs
We first conduct ablation study to demonstrate the contribution of each component in AgentMove for its excellent performance, which are presented in Table 2. We first discuss the impact of three core components individually, as detailed in the top four lines of Table 2. Overall, all components contribute to performance improvement in most cases. However, the performance gains vary across different metrics. For example, while memory design leads to better performance in the Acc@1 in SaoPaulo, the performance in other three metrics are dropped. The effects of the combination of the core components in the last three lines in Table 2. In summary, compared with the base prompt design, the combination of proposed designs introduce 7%-45% performance gain in all the datasets.
Besides, to demonstrate the effects of the World Knowledge Generator (WKG) in exploring new locations, we analyze whether our model explores more potential locations that are not present in the user’s recent contextual stays after incorporating the WKG module. The results are presented in the Table 3. A higher percentage indicates that the model tends to revisit locations from the recent contextual stays, while a lower percentage suggests that the model explores more new locations. The results demonstrate that the WKG successfully encourages the model to explore new locations, which is particularly beneficial for improving performance.
Models FSQ@SaoPaulo ISP@Shanghai
Acc@1 Acc@5 NDCG@5 Acc@1 Acc@5 NDCG@5
base 0.165 0.385 0.277 0.170 0.425 0.298
+STM 0.190 0.315 0.255 0.170 0.445 0.312
+WKG 0.175 0.365 0.269 0.155 0.390 0.276
+CKE 0.175 0.380 0.275 0.175 0.465 0.317
+STM+WKG 0.240 0.390 0.317 0.215 0.455 0.342
AgentMove 0.230 0.415 0.326 0.190 0.450 0.329
vs base 45.45% 7.79% 17.99% 11.76% 5.88% 10.30%
@@
Table 2: Ablation studies of AgentMove. ‘base’ denotes the basic prompts which is similar to the baseline LLM-ZS, ‘+STM’ denotes adding spatial-temporal memory, ‘+WKG’ denotes adding world knowledge generator, ‘+CKE’ denotes adding collective knowledge extractor.
[…]
LLama3-8b w/ WKG 94% 75.6%
w/o WKG 93% 87.8%
LLama3-70b w/ WKG 87.5% 73.2%
w/o WKG 90% 85.4%
Table 3: Effectiveness of word knowledge generator (WKG) for encouraging mobility exploration, which is measured by the location return rates. The location return rate measures the tendency to revisit previously visited locations based on recent contextual stays.
4.4 Geographical Bias and LLM Effects
While LLMs are trained with the online web text which can be geographically bias Manvi et al. (2024) around the world. We investigate the potential geographical bias in LLM based mobility prediction methods and attempt to answer whether AgentMove can alleviate the geographical bias inherent in LLMs to some extent. Experiment results conducted on 12 cities are presented in Figure 4.
In Figure 4(a), we can find significant differences in the accuracy of three LMM-based methods across cities. For instance, cities like Tokyo, Paris, and Sydney generally achieve higher accuracy, while cities like Cape Town and Nairobi see notably lower performance. This suggests the presence of geographical bias in trained LLMs. We also find that proposed AgentMove performs best in most of the cities. Figure 4(b) provides a box-plot test by comparing the Acc@5 of the three LLM-based methods in 12 cities. Results demonstrate that AgentMove not only outperforms the other methods in terms of overall accuracy but also exhibits a smaller range of error. The performance of AgentMove is more consistent across different cities, suggesting a reduced impact of geographical bias with carefully designs in it.
As the core foundation of AgentMove, the capabilities of base LLM play a critical role in its performance. Thus, we evaluate the impact of different LLMs with varying sizes and structures in Figure 5 by using FSQ@Tokyo. Figure 5(a) presents the impact of various 7B LLM with different training data and model structures. The results show that proposed AgentMove performs best adaptability among different LLMs. While LLM-Mob performs stable in all the 7B-LLMs, its performance on Gemma2-9B is far worse than other two methods. We then discuss the detailed impacts of LLM size on AgentMove’s performance across different data. Figure 5(b) reveals that that larger models, particularly Llama3.1-405B, generally deliver significant performance gains for AgentMove compared to smaller models like Llama3-8B across different cities. It is also observed that in Tokyo, Llama3-1-405B performs slightly weaker compared to Llama3-70B. This suggests that while larger models often excel, their effectiveness may vary depending on the unique mobility patterns and characteristics of each city.
@@ 4.4 Geographical Bias and LLM Effects
Figure 4: Geospatial bias analysis of various methods in mobility prediction across 12 cities, where AgentMove outperforms most methods and exhibits lower geospatial bias.
(a) Effects of LLM types on three LLM-based methods
(b) Effects of LLM size on three cities.
@@
Figure 5: The effects of LLM with varying sizes and sources on the prediction performance of three LLM based methods.
5 Related Work
5.1 Mobility Prediction with Deep Learning
Significant efforts have been made in mobility prediction using deep learning models, encompassing research from both sequential-based methods and graph-based methods. Traditional approaches typically employ Markov models Rendle et al. (2010); Cheng et al. (2013) to predict the next visit by learning the transition probabilities between consecutive POIs. In contrast, sequential-based deep learning methods have been proposed to model the high-order movement patterns in trajectory data. These methods can be categorized into two types: recurrent neural networks(RNNs) Kong and Wu (2018); Huang et al. (2019); Yang et al. (2020); Zhao et al. (2020); Feng et al. (2020), and attention mechanismsFeng et al. (2018); Luo et al. (2021); Lin et al. (2021); Cui et al. (2021); Qin et al. (2022); Hong et al. (2023) based works.
Despite their success, these methods primary focus on extracting mobility patterns from an individual perspective, while overlooking the collaborative information available from other users’ trajectories. To address this limitation, recent works Rao et al. (2022); Yang et al. (2022) have explored the use of graph neural network(GNNs) for their ability to model complex relationships. However, all these methods rely on collecting large volumes of private trajectory data. In contrast, our AgentMove leverages the world knowledge and sequential modeling abilities of LLMs to enable the generalized mobility prediction with zero-shot prediction ability.
5.2 Large Language Models and Agents
Due to the powerful language-based generalization and reasoning capabilities Wei et al. (2022a), large language models OpenAI (2022); Touvron et al. (2023) have developed rapidly and have been widely applied in various tasks, including programming Qian et al. (2024) and mathematics Wei et al. (2022a). Recent studies Gurnee and Tegmark (2023); Manvi et al. (2023) have found that LLMs possess a significant amount of geographical knowledge about the world. Additionally, researchers also explore the potential of applying LLMs in spatial-temporal data modelling by directly converting domain-specific tasks into a language-based format, such as time series forecasting Gruver et al. (2024), traffic prediction Li et al. (2024b), trajectory mining Wang et al. (2023); Beneduce et al. (2024), trip recommendation Xie et al. (2024); Li et al. (2024a), traffic signal control Lai et al. (2023); Feng et al. (2024b), comprehensive urban tasks Feng et al. (2024a, b).
These early works highlight the potential of LLMs in spatial-temporal modelling. To effectively utilize the vast knowledge embedded in LLMs and stimulate their reasoning and planning abilities, various prompt techniques Wei et al. (2022b); Kojima et al. (2022); Wang et al. (2022); Yao et al. (2024) have been proposed for solving naive text games and mathematical problems. However, when for more complex real-life and domain-specific tasks, these simple prompt techniques alone are insufficient. Recently, LLM based agents Wang et al. (2024); Xi et al. (2023); Du et al. (2024) are been proposed to address this limitation by equipping LLMs with explicit memory, structured workflows and external tools. In this work, we are the first to design LLM based agent specifically for the mobility prediction task. By incorporating explicit spatial-temporal memory and a workflow for geospatial and social structure mining, we successfully leverage the world knowledge of LLMs and their structured reasoning capabilities for mobility trajectory modelling.
6 Conclusion
In this paper, we propose AgentMove, a systematic agentic framework for generalized human mobility prediction applicable worldwide. We design a spatial-temporal memory module and a collective knowledge extractor to learn both individual mobility patterns and shared mobility pattern among users. Furthermore, we develop a world knowledge generator that utilizes text-based address to understand urban structures in a manner similar to humans. Extensive experiments on trajectories from 12 cities demonstrate the superiority and robustness of AgentMove for mobility prediction.
In the future, we plan to explore more effective ways to extract and leverage the vast world knowledge and common sense of LLMs for mobility modeling and trajectory mining. Additionally, we aim to extend the framework to other trajectory data mining tasks, such as trajectory classification and generation. We believe that LLM-based agents, like AgentMove, hold great potential and adaptability, paving the way for a new paradigm in spatial-temporal modeling alongside deep learning.
7 Limitations
Here, we discuss the potential limitations of the current work and outline directions for future exploration.
Robustness and Hallucination Based on an LLM, the output of AgentMove is not fully controllable. In this work, we define a simple output parser to extract the expected context from the LLM output, but it may occasionally fail. Due to the potential for hallucination in LLMs, the output of AgentMove may include false addresses that do not exist in the real world. While we can define a clear list of valid locations during experiments to verify this, doing so in real-world applications presents significant challenges.
High Cost The high cost of invoking the LLM API limited our experiments to 12 cities with a small test set. This cost will also pose a challenge for large-scale deployment in real-world scenarios. The reliance on LLMs does pose a significant limitation in terms of scalability of our method. With ongoing advancements Liu et al. (2024); Qu et al. (2025) in the development of more efficient and scalable LLM alternatives—such as smaller LLMs, model pruning, and knowledge distillation—we are optimistic about the potential for rapidly decreasing computational costs while improving scalability.
Geospatial Bias Geographical bias has long been a challenging issue in LLMs Manvi et al. (2024). While our proposed AgentMove incorporates specific design elements to mitigate some of these biases, it cannot completely eliminate them due to inherent limitations in LLMs. However, we believe that our current work represents a significant step forward in addressing geospatial bias in mobility prediction challenges. One promising direction for further reducing geographical bias could be the integration of more external knowledge during inference, and we are actively exploring this avenue in our future work.
8 Ethics Statement
All trajectory data used in the experiments come from publicly available open-source datasets Yang et al. (2016); Feng et al. (2019). We do not attempt to extract any personal information from these datasets.